# Discrete Time Crystal Phase as a Resource for Quantum Enhanced Sensing

### v2

@@ Abstract
Figure 2: (a) The QFI $\mathcal{F}_{Q}$ versus $\omega$ in stroboscopic times in a system of size $L{=}30$. The onset of the phase transition is determined by $\omega{=}\omega_{\max}$ (dashed line), the point where QFI peaks in different $n$’s. (b) Dynamical growth of the QFI when a system of size $L{=}12$ is deeply in DTC phase (for $\omega{=}10^{-4}\pi/2$), in the non-DTC phase (for $\omega{=}10^{-1}\pi/2$), and at the transition point ($\omega{=}\omega_{\max}$). (c) and (d) QFI versus $\omega$ in systems with various $L$’s after $n{=}2$ and $n{=}100$, respectively. Inset: $\omega_{\max}$ versus $L$ at $n{=}2$ and $n{=}100$. The numerical simulation (NS) are well described by the fitting function $\omega_{\max}{\propto}L^{-1}$. (e) The values of the QFI after $n=2$ in DTC phase (for $\omega{=}10^{-4}\pi/2$) and at transition points ($\omega{=}\omega_{\max}$) versus $L$. The numerical simulation (NS) is well-mapped by a function as $\mathcal{F}_{Q}{\propto}L^{\beta}$ (solid lines) with $\beta{>}3$. (f) The finite-size scaling analysis obtained for the curves in panel (c). The best data collapse is obtained for the reported critical parameters ($\omega_{c},\zeta,\nu$). Here, $Q$ determines the quality of the data collapse with $Q{=}1$ for the optimal data collapse Melchert (2009); Sorge (2015).
@@
DTC sensor.– To investigate the sensing capability of our DTC probe for sensing $\omega$, in Fig. 2(a), we plot QFI $\mathcal{F}_{Q}$ as a function of $\omega$ at different stroboscopic times $n{\in}\{2,10,\cdots,100\}$, in a chain of length $L{=}30$ and $\varepsilon{=}0.01$. Several interesting features can be observed. First, the QFI shows distinct behaviors in each phase. While in the DTC phase, the $\mathcal{F}_{Q}$ becomes a plateau whose value depends on $n$, in the non-DTC region it shows nontrivial and fast oscillations. Second, by approaching the transition point, denoted by $\omega_{\max}$ (dashed line), the QFI indeed shows a clear peak at all stroboscopic times. Note that $\omega_{\max}$ in both Fig. 1(b) and Fig. 2(a) are exactly the same. To understand the dynamical growth of the QFI, in Fig. 2(b), we plot $\mathcal{F}_{Q}$ over thousands of driving cycle $n$ in a systems of size $L{=}12$ and $\varepsilon{=}0.01$ at different $\omega$’s. As the figure shows, when the probe is tuned to work deeply in either DTC phase (for $\omega{=}10^{-4}\pi/2$) or the non-DTC region (for $\omega{=}10^{-1}\pi/2$), one obtains $\mathcal{F}_{Q}{\propto}n^{2}$. However, in the transition point, $\omega{=}\omega_{\max}$ the QFI in the early times $n{\in}[1,\cdots,80]$ dramatically increases as $\mathcal{F}_{Q}{\propto}n^{3}$, and then follows $\mathcal{F}_{Q}{\propto}n^{2}$ in the larger times. To identify the effect of size on quality of sensing, we analyze the QFI at various sizes $L{=}12,\cdots,32$ and also different cycles $n$. In Figs. 2(c) and (d), we plot the obtained $\mathcal{F}_{Q}$ as a function of $\omega$ after $n{=}2$ and $n{=}100$ cycling periods, respectively, for various $L$ and fixed $\varepsilon{=}0.01$. The finite-size effect is obvious in both DTC phase and transition point, namely, the points where QFI peaks. By enlarging the chain, the peaks of the QFI smoothly skew towards smaller values of $\omega$ as can be seen in the inset of Fig. 2(d). The obtained $\omega_{\max}$ at different $n$’s are well-mapped with function $\omega_{\max}{\propto}L^{-1}$, indicating that in the thermodynamic limit $L{\rightarrow}\infty$ the transition happens at infinitesimal deviation $\omega$. In the non-DTC region, the oscillatory behavior of the QFI, in particular in larger times ($n{=}100$) prevents us from investigating the scaling behaviors in this phase. In Fig. 2(e), we present the QFI at $n{=}2$ as a function of $L$ at $\omega{=}10^{-4}\pi/2$, namely deep inside the DTC phase, and also at the corresponding transition points $\omega{=}\omega_{\max}$. The numerical results can be properly mapped with a fitting function as $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}3.125$ and $\beta{=}3.13$ in the DTC phase and at the transition point, respectively. Based on these, one can suggest the following ansatz for the QFI
@@
where throughout the DTC phase one has $\alpha{\simeq}2$ and $\beta{\simeq}3$. We highlight this as the main result of this Letter showing that our DTC probe achieves quantum-enhanced sensitivity. It is worth emphasizing that in classical probes one at best achieves $\beta{=}1$. Exploiting quantum features may enhance the precision to $\beta{=}2$, conventionally known as the Heisenberg limit. The ultimate obtainable precision in $k$-body interacting system, however, is given by $\beta{=}2k$, which becomes $\beta{=}4$ in our case Boixo et al. (2007). Thus, while our DTC probe beats the conventional Heisenberg limit, it remains below the ultimate precision bound.
@@
To characterize the observed phase transition, we assume a continuous second-order ansatz for the QFI as $\mathcal{F}_{Q}{=}L^{\zeta/\nu}\mathcal{G}(L^{1/\nu}(\omega{-}\omega_{c}))$ where $\zeta$ and $\nu$ are critical exponents, $\omega_{c}$ is the critical point and $\mathcal{G}$ is an arbitrary function. If this ansatz is correct, one expects to obtain data collapse of various size systems when $L^{-\zeta/\nu}\mathcal{F}_{Q}$ is plotted versus $L^{1/\nu}(\omega{-}\omega_{c})$. Indeed, as shown in Fig. 2(f), tuning the parameters to $(\omega_{c},\zeta,\nu){\simeq}(0.00026,2.9569,0.9488)$, optimized using Python package PYFSSA Melchert (2009); Sorge (2015), results in an almost perfect data collapse for curves in Fig. 2(c). This indicates that the DTC phase transition is indeed of the second-order type.
Figure 3: (a) The QFI versus $\omega$ in a chain of size $L{=}20$ after $n{=}10$ period cycles for different $\varepsilon$. The dashed line determines the onset of the phase transition. (b) The QFI versus $\varepsilon$ for two values of $\omega$, obtained after $n{=}100$ cycling periods in a system of size $L{=}12$.
Imperfection effect.– Previously we analytically proved that our DTC is robust against uniform imperfection $\varepsilon$ in the driving pulse when $\omega{=}0$ (see SM). In the case of nonzero $\omega$, the situation becomes even more interesting. In Fig. 3(a), we plot the QFI versus $\omega$ after $n{=}10$ for a chain of size $L{=}20$ under driving pulse with various imperfections $\varepsilon$. While the qualitative behavior of the probe in the DTC phase is not affected by imperfection, increasing $\varepsilon$ may enhance the QFI. To assess the performance in a wider range of the imperfection, in Fig. 3(b) we report $\mathcal{F}_{Q}$ as a function of $\varepsilon$ for $\omega{=}10^{-4}\pi/2$, and $\omega{=}\omega_{\max}$. The results are obtained after $n{=}100$ in a system of size $L{=}12$. This can be understood as imperfect rotating pulses through involving a larger sector of the Hilbert space in the dynamics of the system imprints more information about $\omega$ into the quantum state. The enhancement in the DTC phase, where the system is supposed to be strongly localized, has remarkably stronger effect than at the transition point that already has features of both thermalization and localization. This interesting result is in sharp contrast with the usual sensors where the imperfections deteriorate the sensing power.
@@
Melting transition of the DTC.– Having elucidated the sharp second-order phase transition controlled by $\omega$, we now explore the melting of the DTC by increasing the imperfection $\varepsilon$. In Fig. 4(a), we plot the revival fidelity $F(nT)$ after $n{=}100$ period cycles as a function of $\omega$ and $\varepsilon$, obtained for system of size $L{=}12$. Indeed the phase diagram is fully described by both $\omega$ and $\varepsilon$. To diagnose the transition driven by the imperfection $\varepsilon$, we use averaged entanglement entropy $\langle S_{EE}\rangle$ and averaged diagonal entropy $\langle S_{DE}\rangle$ obtained for all Floquet states of $U_{F}(\Omega,\varepsilon){=}\sum_{k=1}^{2^{L}}e^{-i\phi_{k}}|\phi_{k}\rangle\langle\phi_{k}|$. For a given Floquet state $|\phi_{k}\rangle$, the reduced density matrix $\rho_{L/2}^{(k)}=\mathrm{Tr}_{L/2}|\phi_{k}\rangle\langle\phi_{k}|$ can be obtained by tracing out $L/2$ spins in the right side of the chain. Therefore, the entanglement entropy between the half-systems is $S_{EE}^{(k)}{=}{-}\mathrm{Tr}[\rho_{L/2}^{(k)}\ln(\rho_{L/2}^{(k)})]$ with an average as $\langle S_{EE}\rangle{=}\sum_{k=1}^{2^{L}}S_{EE}^{(k)}/2^{L}$. Replacing $\rho_{L/2}^{(k)}$ by decohered density matrix $\varrho_{L/2}^{(k)}$ which only contain the diagonal elements of $\rho_{L/2}^{(k)}$, results in diagonal entropy as $S_{DE}^{(k)}{=}{-}\mathrm{Tr}[\varrho_{L/2}^{(k)}\ln(\varrho_{L/2}^{(k)})]$ and its average $\langle S_{DE}\rangle{=}\sum_{k=1}^{2^{L}}S_{DE}^{(k)}/2^{L}$. This quantity has recently been proposed for emulating the thermodynamic behavior in many-body localization contexts Yousefjani and Bayat (2023); Yousefjani et al. (2023b). In the DTC phase, each Floquet state is a maximally entangled Greenberger-Horne-Zeilinger (GHZ) of two computational basis states. For instance, for negligible values of the $\varepsilon$, one approximately has $|\phi_{1}\rangle{\cong}\frac{1}{\sqrt{2}}(|\mathbf{0}\rangle+|{\mathbf{2^{L}-1}}\rangle)$ and $|\phi_{2^{L}}\rangle{\cong}\frac{1}{\sqrt{2}}(|\mathbf{0}\rangle-|\mathbf{2^{L}-1}\rangle)$ with the corresponding eigenvalues as $\phi_{1}{\cong}E_{\mathbf{1}}$ and $\phi_{2^{L}}{\cong}E_{\mathbf{1}}\pm\pi$. This so-called $\pi$-pairs of the Floquet states results in $S_{EE}^{(k)}{=}S_{DE}^{(k)}{=}\ln 2$ in deep DTC phase. By increasing $\varepsilon$ both entanglement and diagonal entropy grow and peak at $\varepsilon{=}0.5$, see Fig. 4(b) and also SM. To extract the critical properties of a transition, one can establish finite-size scaling analysis to predict the behavior of the system when its size is changing. In the inset of Fig. 4(b), we depict the best collapse of the corresponding curves obtained for reported $(\varepsilon_{c},\nu,\zeta)$. By increasing the system size $\omega_{\max}$ decreases and thus the extension of the DTC phase becomes smaller. Therefore, for finite-size scaling analysis we select the lengths such that for the given $\omega$ they are all within the DTC phase when $\varepsilon{\simeq}0$. Note that the obtained $\varepsilon_{c}$ from finite-size scaling of both $\langle S_{EE}\rangle$ and $\langle S_{DE}\rangle$, shown as markers on panel (a), are very close and determine the phase boundary between the DTC and the non-DTC phase.
@@
Non-DTC region.– By increasing $\omega$, one observes non trivial oscillations in the behvior of fidelity $F(2nT)$, see Fig. 1(a), which signals the lost of the DTC order. As the system size increases the period of these oscillations increases, hinting that the system thermalizes in the thermodynamic limit $L{\rightarrow}\infty$. As shown in the SM, the entanglement entropy $\langle S_{EE}\rangle$ and the diagonal entropy $\langle S_{DE}\rangle$ asymptotically approach their corresponding Page entropy Page (1993); Torres-Herrera and Santos (2017); Torres-Herrera et al. (2016), expected for ergodic thermal phases, as the system size increases. This provides further evidence that the non-DTC phase becomes an ergodic thermal phase in the thermodynamic limit.

### v3

@@ 5 DTC sensor
Figure 2: (a) The QFI $\mathcal{F}_{Q}$ versus $\omega$ in stroboscopic times in a system of size $L{=}30$. The onset of the phase transition is determined by $\omega{=}\omega_{\max}$ (dashed line), the point where QFI peaks in different $n$’s. (b) Dynamical growth of the QFI when a system of size $L{=}12$ is deeply in DTC phase (for $\omega{=}10^{-4}\pi/2$), in the non-DTC phase (for $\omega{=}10^{-1}\pi/2$), and at the transition point ($\omega{=}\omega_{\max}$). (c) and (d) QFI versus $\omega$ in systems with various $L$’s after $n{=}2$ and $n{=}100$, respectively. Inset: $\omega_{\max}$ versus $L$ at $n{=}2$ and $n{=}100$. The numerical simulation (NS) are well described by the fitting function $\omega_{\max}{\propto}L^{-1}$. (e) The values of the QFI after $n=2$ in DTC phase (for $\omega{=}10^{-4}\pi/2$) and at transition points ($\omega{=}\omega_{\max}$) versus $L$. The numerical simulation (NS) is well-mapped by a function as $\mathcal{F}_{Q}{\propto}L^{\beta}$ (solid lines) with $\beta{>}3$. (f) The finite-size scaling analysis obtained for the curves in panel (c). The best data collapse is obtained for the reported critical parameters ($\omega_{c},\zeta,\nu$). Here, $Q$ determines the quality of the data collapse with $Q{=}1$ for the optimal data collapse [105, 106].
@@
To investigate the sensing capability of our DTC probe for sensing $\omega$, in Fig. 2(a), we plot QFI $\mathcal{F}_{Q}$ as a function of $\omega$ at different stroboscopic times $n{\in}\{2,10,\cdots,100\}$, in a chain of length $L{=}30$ and $\varepsilon{=}0.01$. Several interesting features can be observed. First, the QFI shows distinct behaviors in each phase. While in the DTC phase, the $\mathcal{F}_{Q}$ becomes a plateau whose value depends on $n$, in the non-DTC region it shows nontrivial and fast oscillations. Second, by approaching the transition point, denoted by $\omega_{\max}$ (dashed line), the QFI indeed shows a clear peak at all stroboscopic times. Note that $\omega_{\max}$ in both Fig. 1(b) and Fig. 2(a) are exactly the same. To understand the dynamical growth of the QFI, in Fig. 2(b), we plot $\mathcal{F}_{Q}$ over thousands of driving cycle $n$ in a systems of size $L{=}12$ and $\varepsilon{=}0.01$ at different $\omega$’s. Clearly, when the probe is tuned to work deeply in either DTC phase (for $\omega{=}10^{-4}\pi/2$) or the non-DTC region (for $\omega{=}10^{-1}\pi/2$), one obtains $\mathcal{F}_{Q}{\propto}n^{2}$. However, in the transition point, $\omega{=}\omega_{\max}$ the QFI in the early times $n{\in}[1,\cdots,80]$ dramatically increases as $\mathcal{F}_{Q}{\propto}n^{3}$, and then follows $\mathcal{F}_{Q}{\propto}n^{2}$ in the larger times. To identify the effect of size on quality of sensing, we analyze the QFI at various sizes $L{=}12,\cdots,32$ and also different cycles $n$. In Figs. 2(c) and (d), we plot $\mathcal{F}_{Q}$ as a function of $\omega$ after $n{=}2$ and $n{=}100$ cycling periods, respectively, for various $L$ and fixed $\varepsilon{=}0.01$. The finite-size effect is obvious in both DTC phase and transition point. By enlarging the chain, the peaks of the QFI smoothly skew towards smaller $\omega$, see the inset of Fig. 2(d). The obtained $\omega_{\max}$ at different $n$’s are well-mapped with function $\omega_{\max}{\propto}L^{-1}$, indicating that in the thermodynamic limit $L{\rightarrow}\infty$ the transition happens at infinitesimal deviation $\omega$. In the non-DTC region, QFI oscillations, especially over extended periods ($n{=}100$), hinder scaling behavior investigation. In Fig. 2(e), we present the QFI at $n{=}2$ as a function of $L$ at $\omega{=}10^{-4}\pi/2$, namely deep inside the DTC phase, and also at the corresponding transition points $\omega{=}\omega_{\max}$. The numerical results can be properly mapped with a fitting function as $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}3.125$ and $\beta{=}3.13$ in the DTC phase and at the transition point, respectively. Based on these, one can suggest the following ansatz for the QFI
@@
where throughout the DTC phase one has $\alpha{\simeq}2$ and $\beta{\simeq}3$. We highlight this as the main result of this Letter showing that our DTC probe achieves quantum-enhanced sensitivity. It is worth emphasizing that in classical probes one at best achieves $\beta{=}1$. Exploiting quantum features may enhance the precision to $\beta{=}2$, known as the Heisenberg limit. The ultimate precision in $k$-body interacting systems is $\beta{=}2k$, which equals $\beta{=}4$ in our case [94].
@@
To study the observed phase transition, we start with a continuous second-order ansatz for the QFI as $\mathcal{F}_{Q}{=}L^{\zeta/\nu}\mathcal{G}(L^{1/\nu}(\omega{-}\omega_{c}))$ where $\zeta$ and $\nu$ are critical exponents, $\omega_{c}$ is the critical point and $\mathcal{G}$ is an arbitrary function. If this ansatz is correct, one expects to obtain data collapse of various size systems when $L^{-\zeta/\nu}\mathcal{F}_{Q}$ is plotted versus $L^{1/\nu}(\omega{-}\omega_{c})$. Indeed, as shown in Fig. 2(f), tuning the parameters to $(\omega_{c},\zeta,\nu){\simeq}(0.00026,2.9569,0.9488)$, optimized using Python package PYFSSA [105, 106], results in an almost perfect data collapse for curves in Fig. 2(c). This indicates that the DTC phase transition is indeed of the second-order type.
6 Melting transition of the DTC
@@
Having elucidated the sharp second-order phase transition controlled by $\omega$, we now explore the melting of the DTC by increasing the imperfection $\varepsilon$. In Fig. 3(a), we plot the revival fidelity $F(nT)$ after $n{=}100$ as a function of $\omega$ and $\varepsilon$, for system of size $L{=}12$. Indeed the phase diagram is fully described by $\omega$ and $\varepsilon$. To diagnose the transition driven by the imperfection $\varepsilon$, we use averaged entanglement entropy $\langle S_{EE}\rangle$ and averaged diagonal entropy $\langle S_{DE}\rangle$ obtained for all Floquet states of $U_{F}(\Omega,\varepsilon){=}\sum_{k=1}^{2^{L}}e^{-i\phi_{k}}|\phi_{k}\rangle\langle\phi_{k}|$. For a given Floquet state $|\phi_{k}\rangle$, the reduced density matrix $\rho_{L/2}^{(k)}=\mathrm{Tr}_{L/2}|\phi_{k}\rangle\langle\phi_{k}|$ can be obtained by tracing out $L/2$ spins in the right side of the chain. Therefore, the entanglement entropy between the half-systems is $S_{EE}^{(k)}{=}{-}\mathrm{Tr}[\rho_{L/2}^{(k)}\ln(\rho_{L/2}^{(k)})]$ with an average as $\langle S_{EE}\rangle{=}\sum_{k=1}^{2^{L}}S_{EE}^{(k)}/2^{L}$. Replacing $\rho_{L/2}^{(k)}$ by decohered density matrix $\varrho_{L/2}^{(k)}$ which only contain the diagonal elements of $\rho_{L/2}^{(k)}$, results in diagonal entropy as $S_{DE}^{(k)}{=}{-}\mathrm{Tr}[\varrho_{L/2}^{(k)}\ln(\varrho_{L/2}^{(k)})]$ and its average $\langle S_{DE}\rangle{=}\sum_{k=1}^{2^{L}}S_{DE}^{(k)}/2^{L}$.
@@
This quantity has recently been proposed for emulating the thermodynamic behavior in many-body localization contexts [99, 101]. In the DTC phase, each Floquet state is a maximally entangled Greenberger-Horne-Zeilinger (GHZ) of two computational basis states. For instance, for negligible $\varepsilon$, one has $|\phi_{1}\rangle{\cong}\frac{1}{\sqrt{2}}(|\mathbf{0}\rangle{+}|{\mathbf{2^{L}-1}}\rangle)$ and $|\phi_{2^{L}}\rangle{\cong}\frac{1}{\sqrt{2}}(|\mathbf{0}\rangle{-}|\mathbf{2^{L}-1}\rangle)$ with the corresponding eigenvalues as $\phi_{1}{\cong}E_{\mathbf{1}}$ and $\phi_{2^{L}}{\cong}E_{\mathbf{1}}\pm\pi$. This so-called $\pi$-pairs of the Floquet states results in $S_{EE}^{(k)}{=}S_{DE}^{(k)}{=}\ln 2$ in deep DTC phase. By increasing $\varepsilon$ both entanglement and diagonal entropy grow and peak at $\varepsilon{=}0.5$, see Fig. 3(b) and appendix C. To characterize critical properties, finite-size scaling analysis needs to be established. In the inset of Fig. 3(b), we depict the best collapse of the corresponding curves obtained for reported $(\varepsilon_{c},\nu,\zeta)$. By increasing length $\omega_{\max}$ decreases and thus the extension of the DTC phase becomes smaller. Therefore, for finite-size scaling analysis we select the lengths such that for the given $\omega$ they are all within the DTC phase when $\varepsilon{\simeq}0$. Note that the obtained $\varepsilon_{c}$ from finite-size scaling of both $\langle S_{EE}\rangle$ and $\langle S_{DE}\rangle$, shown as markers on panel (a), are very close and determine the phase boundary between the DTC and the non-DTC phase. Note that, while strong imperfection, i.e. large $\varepsilon$, melts the DTC, small imperfections may indeed enhance the sensitivity of our probe, see appendix C.
Figure 4: (a) Quantum circuit for implementing our DTC on a digital quantum simulators. The operator $U_{F}(\Omega,\varepsilon)$ is realized through a combination of CPHASE gates and $ZZ(j\Omega)$ and single-qubit rotations $X(\Phi)$. (b) Time-averaged CFI, ¯$\overline{\mathcal{F}}_{C}$, over $n{=}10$ as a function of $L$ in DTC phase, $\omega{=}10^{-4}\pi/2$, and at the transition points, $\omega{=}\omega_{\max}$. Inset is ¯$\overline{\mathcal{F}}_{C}$, and time-averaged QFI, ¯$\overline{\mathcal{F}}_{Q}$, as a function of $\omega$ for $L{=}20$.
7 Non-DTC region
By increasing the divination $\omega$, the perfect and stable revivals of the fidelity in stroboscopic times, namely $F(2nT){=}1$ for $\omega{\leq}\omega_{\max}$, are replaced by nontrivial oscillations for $\omega{>}\omega_{\max}$. This hints one enters a non-DTC region. In this section, we aim to characterize the nature of this region. Our results for the revival fidelity $F(2nT)$ as a function of $n$ for systems of different sizes that are tuned to work in the non-DTC region, namely for $\omega{\simeq}0.15$, have been illustrated in Fig. 5(a). By enlarging the system size, the period of these incommensurate fluctuations increases. This implies that, in systems with enough large sizes, these oscillations practically vanish in a reasonable time window, signaling the thermalization of the system. This observation receives more support from our static study based on the entanglement entropy and diagonal entropy. In a thermal system, the Floquet states $\{|\phi_{k}\rangle\}$ are expected to behave as a typical random pure state, therefore their entanglement entropy is predicted to follow the Page entropy $\langle S^{P}_{EE}\rangle{\simeq}(L\ln(2)-1)/2$ for enough large $L$’s [107, 108, 109]. In this case, the average entanglement entropy should already captured its maximum and the variations of $\varepsilon$ may not considerably affect $\langle S^{P}_{EE}\rangle$. Our numerical results in Fig. 5(b) support this prediction for $\omega{=}\frac{\pi}{4}$, namely deep inside the thermal phase. The results for systems of size $L{=}8$ and $L{=}12$ capture the Page entropy $\langle S_{EE}\rangle{\simeq}\langle S^{P}_{EE}\rangle$ (depicted by colored dashed lines) with slight changes in terms of $\varepsilon$. Regarding the diagonal entropy, typical random pure states are expected to follow $\langle S^{P}_{DE}\rangle{\simeq}\ln(0.48\times 2^{L/2})+\ln(2)$. The presented results in Fig. 5(c) confirm this behavior. In this panel $\langle S^{P}_{DE}\rangle$ is represented by colored dashed lines.
Figure 5: (a) the stroboscopic dynamics of the revival fidelity $F(2nT)$ over $100$ period cycles when systems with different sizes evolve in non-DTC region with $\omega{\simeq}0.15$ and $\varepsilon{=}0.01$. (b) the averaged entanglement entropy $\langle S_{EE}\rangle$ and (c) the averaged diagonal entropy $\langle S_{DE}\rangle$ as a function of pulse imperfection $\varepsilon$ obtained for systems of different sizes and $\omega{=}\pi/4$. Dashed lines in (b) and (c) determine the corresponding $\langle S^{P}_{EE}\rangle$ and $\langle S^{P}_{DE}\rangle$, respectively.
8 Experimental realization
We propose superconducting quantum simulators, similar to Ref. [35], for realizing our DTC. Rewriting the Floquet operator $U_{F}(\Omega,\varepsilon)=e^{-iH_{P}}e^{-i\Omega H_{I}}$ as
$U_{F}(\Omega,\varepsilon)=\Big{(}\otimes_{j=1}^{L}e^{-i\Phi\sigma^{x}_{j}}\Big{)}\Big{(}\otimes_{j=1}^{L-1}e^{-ij\Omega\sigma^{z}_{j}\otimes\sigma^{z}_{j+1}}\Big{)},$ (5)
clearly show that $U_{F}$ is an exact combinations of controlled-phase (CPHASE) gate $ZZ(j\Omega){=}e^{-ij\Omega\sigma^{z}_{j}\otimes\sigma^{z}_{j+1}}$ followed by single-qubit rotation $X(\Phi){=}e^{-i\Phi\sigma^{x}_{j}}$. The proposed protocol, see Fig. 4 (a), for implementing our DTC includes three parts. After initializing the superconducting qubits in an arbitrary computational basis $|\mathbf{z}\rangle$, the Floquet operator $U_{F}$ is implemented identically $n$ times. Then at each cycle period the local polarization of each qubit $\langle\sigma^{z}_{j}\rangle$ is measured to track the overlap of the evolved state with $|\mathbf{z}\rangle$. The newly developed tunable CPHASE gates allow us to engineer the gradient interaction between qubits to fulfill the requirement of our DTC. The technical details of the implementation and calibration of this two-qubit gate can be found in Ref. [35]. In an array of $L$ qubits, implementing a single-period evolution $U_{F}$ requires $L{-}1$ two-qubit gates and $L$ single-qubit rotations. The standard thermal relaxation time, $T_{1}$, and dephasing time, $T_{2}^{*}$, that characterize the typical operational time interval of the superconducting simulator have been reported as $T_{1}{\simeq}16$ $\mu$s and $T_{2}^{*}{\simeq}6$ $\mu$ [35]. Therefore, setting the CPHASE gate duration to ${\sim}20{+}20$ ns and the single-qubit gate duration to ${\sim}20$ ns allows us to excuse $n{=}T_{2}^{*}/0.06{\simeq}100$ cycling period in the operational interval of the circuit. Our analysis shows that a simple configuration measurement described by projective operators as $\{\Pi_{\mathbf{z}}{=}|\mathbf{z}\rangle\langle\mathbf{z}|\}$ can saturate the quantum Cramér-Rao bound. Here, $\{|\mathbf{z}\rangle\}$ represents the $2^{L}$ different computational basis. For $p_{\mathbf{z}}{=}|\langle\mathbf{z}|\psi_{n}\rangle|^{2}$ (with $|\psi_{n}\rangle{=}U_{F}^{n}(\Omega,\varepsilon)|\psi_{0}\rangle$) as the probability of finding the spins in the computational basis $|\mathbf{z}\rangle$, one can calculate the CFI. As it is clear from the inset of Fig. 4 (b), the time-averaged CFI, ¯$\overline{\mathcal{F}}_{C}{=}\frac{1}{N}\sum_{n=1}^{N}\mathcal{F}_{C}(nT)$, highly resembles the time-averaged QFI, ¯$\overline{\mathcal{F}}_{Q}{=}\frac{1}{N}\sum_{n=1}^{N}\mathcal{F}_{Q}(nT)$. It shows two distinct behaviors in both DTC phase and non-DTC region and provides a sharp peak near the transition point $\omega_{\max}$. To extract the scaling behavior of the CFI, one can plot ¯$\overline{\mathcal{F}}_{C}$ as a function of $L$ at different $\omega$. Our results presented in Fig. 4 (b) show that the numerical simulations are properly mapped with a fitting function as ¯$\overline{\mathcal{F}}_{C}{\propto}L^{\beta}$ with $\beta{=}3.158$ and $\beta{=}3.20$ for DTC phase and the transition point. This numerical simulation shows that the simple configuration measurement in our DTC sensor can saturate the quantum Cramér-Rao bound and capture the quantum-enhanced sensitivity.
9 Conclusion